Perry wants to make a 80% alcohol solution. He has already poured 8 gallons of a 85% alcohol solution into a beaker. How many gallons of a 70% alcohol solution must be added to this to create the desired mix?
Answer by checkley77(12844)

.85*8+.70x=.80(8+x)
6.8+.7x=6.4+.8x
.7x-.8x=6.4-6.8
-.1x=-.4
x=-.4/-.1
x=4 gallons of 70% alcohol is used.
Proof:
85*8+.7*4=.8(8+4)
6.8+2.8=.8*12
9.6=9.6